# exampledefaultouterlinewidth=0pt,innerlinewidth=0pt, roundcorner=5pt,backgroundcolor=myblue, leftmargin=0.2cm,rightmargin=0.2cm

### v1

@@ 8.2 Least squares setting
In the following we consider three cases, a strongly convex risk in which the spectrum of the eigenvalues is bounded from below (section 8.2.1). A case in which the spectrum is not bounded from below as $d\to\infty$, but the number of eigenvalues below some fixed threshold is $o(d)$ (section 8.2.2). Finally, power law spectrum supported on $[0,1]$ with $d\to\infty$ (section 8.2.3).
8.2.1 Proofs for case of fixed $d$
@@
On the hand following Lemma 8.3, $\frac{1}{4}\mathscr{D}^{2}(0)(1+\frac{\gamma_{0}}{2d}\operatorname{Tr}(K))\leq\|r\|_{1}$. Therefore, $\|r\|_{1}\asymp\frac{1}{4}\mathscr{D}^{2}(0)$.
Next, rewriting the $\gamma_{t}$ equation in Table LABEL:table:learning_rate_expressions (AdaGrad-Norm for least squares) in terms of $g(u)$ (Lemma 8.1), we obtain
@@
Remark 8.1.
We note that, on the Least square problem $\hat{L}(f)=\hat{\mu}(f)=1$, therefore, the bound in Proposition 8.1 yields $\frac{\eta^{2}}{\frac{b}{\eta}+\frac{1}{2(1-\zeta)}\frac{1}{d}\operatorname{Tr}(K)\mathscr{D}^{2}(0)}$.
@@
8.2.2 Vanishingly few eigenvalues near 0 as $d\to\infty$
@@
with $\mathscr{R}_{2}(s)={\frac{1}{2d}}\sum_{i:\lambda_{i}\leq\alpha_{s}}\lambda_{i}\mathscr{D}_{i}^{2}(s)$. Next, choosing $\alpha_{s}=\gamma_{s}\frac{1}{d}\operatorname{Tr}(K^{2})$, we show that the last term is of order $o_{d}(1)$. By Lemma 8.2 $\forall i$, $\mathscr{D}^{2}_{i}(t)\leq\max\left({\gamma_{t_{1}}\mathscr{R}(t_{1})},\mathscr{D}_{i}^{2}(0)\right)={c_{0}}$ where the bound ${c_{0}}$ comes from the assumption $\langle{X^{\star},\omega_{i}}\rangle=O(d^{-1/2})$ and the initialization $X_{0}=0$. Therefore,
@@
8.2.3 Power law behavior at $d\to\infty$
@@
In this section, we use the convolutional Volterra structure of the risk (Eq. (70)) to derive non-asymptotic bounds on the risk, which will be useful in Section 8.2.3 to derive the asymptotic behavior of the risk and the learning rate under power law assumption on the spectrum of the covariance matrix and the discrepancy from the target at initialization.
Lemma 8.3.
@@
where the third transition is by Lemma 8.4, with $C_{1}=\frac{\mathcal{K}(0)}{\mathcal{K}(T)(2\epsilon+1)}+1,$ which then completes the proof. ∎
Lemma 8.4 (Lemma IV.4.7 in [3]).
@@
Corollary 8.1 apply Lemma 8.3 in the setting for which $F$, and $\mathcal{K}$ has a power law behavior asymptotically. It shows that the risk will then be dominated by $F$ only. Corollary 8.2 shows the behavior of the learning rate in this setting. Finally, Lemma 8.5 shows that Assumption 8 is a consequence of a power law spectrum near zero on the eigenvalues of the covariance matrix and a power law assumption on the projected discrepancy at initialization.
Corollary 8.1.
@@
Define $g(u)=\gamma_{\Gamma^{-1}(u)}$ and $r(u)=\mathscr{R}(\Gamma^{-1}(u))$ and observe that $g(u)$ is a decreasing function. Then, from the upper bound in Lemma 8.3, we have
@@
Combining this upper bound with the lower bound from Lemma 8.3 and that $\kappa_{2}>1$, we conclude that $r(u)\asymp F(u)$ and $\mathscr{R}(t)\asymp F(\Gamma(t)).$
@@
Using the notations $g(u)$ and $r(u)$ defined above along with the change of variable $u=\Gamma(t)$, we get $\int_{0}^{t}\mathscr{R}(s)\mathop{}\!\mathrm{d}s=\int_{0}^{u}\frac{r(v)}{g(v)}\mathop{}\!\mathrm{d}v$. Combining this with Corollary 8.1 and the formula for $\gamma_{t}$ we get
@@
In the case of $\kappa_{1}<1$, we have $\sqrt{I(u)}-\sqrt{I(0)}\asymp(1+v)^{1-\kappa_{1}}$ so, for $u$ sufficiently large, $g(u)\asymp(1+u)^{\kappa_{1}-1}$. To recover the asymptotic for $\gamma_{t}$, we observe that $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}u}\Gamma^{-1}(u)=\frac{1}{g(u)}\asymp(1+u)^{1-\kappa_{1}}$. Integrating both sides and changing back to $t$ variables, we get $t\asymp(1+\Gamma(t))^{2-\kappa_{1}}$ (or equivalently $1+\Gamma(t)\asymp t^{1/(2-\kappa_{1})}$). Finally, plugging this into the formula for $\gamma_{t}$ and applying Corollary 8.1, we get
@@
In the case of $\kappa_{1}=1$, we follow a similar procedure as for $\kappa_{1}<1$ to show that $t\asymp\Gamma(t)\log(\Gamma(t))$ for sufficiently large $t$. This implies $\Gamma(t)\asymp t/\log(t)$ which gives the desired result after integration. The decay rate of the risk is then immediate using Corollary 8.1. ∎
Lemma 8.5.
@@
The proof is an immediate application of Corollary 8.2 with, $\kappa_{1}=2-\beta-\delta$ as implied by Lemma 8.5. ∎
Remark 8.2.
@@
9 Polyak Stepsize
The distance to optimality of SGD is measured say by $D^{2}(X)=\|X-X^{\star}\|^{2}$. Let us consider the deterministic equivalent for the distance to optimality $\mathscr{D}^{2}(t)$ in (9). Fixing $T>0$ and any $\varepsilon\in(0,1/2)$, we have by Theorem 2.1 (see also corollary 6.1 which show concentration for large class of statistics) that $\sup_{0\leq t\leq T}|\|X_{\lfloor td\rfloor}-X^{\star}\|^{2}-\mathscr{D}^{2}(t)|\leq d^{-\varepsilon}$, w.o.p. In this way, if we want to guarantee that the distance to optimality of SGD decreases, we need $\mathop{}\!\mathrm{d}\mathscr{D}^{2}(t)<0$ with the maximum decrease being $\min_{\gamma_{t}}\mathop{}\!\mathrm{d}\mathscr{D}^{2}(t)$.
@@ 9 Polyak Stepsize
Figure 5: Convergence in Exact Line Search on a noiseless least squares problem. The plot on the left illustrates the convergence of the risk function, while the center and right plots depict the convergence of the quotient $\frac{\mathscr{D}_{\lambda_{2}}(t)}{\mathscr{D}_{\lambda_{1}}(t)}$ and the learning rate $\gamma_{t}$, respectively. Further details and formulas for the limiting behavior can be found in the Appendix 10.2. See Appendix 12 for simulation details.
10 Line Search
10.1 General Line Search
@@
10.2 Line Search on least squares
@@ 11 Examples
11.1 Binary logistic regression
@@
We note that, the logistic regression is $(\mu,\theta)$–RSI with $\mu=\tfrac{1}{\ell e^{\sqrt{4\theta}}}$ see section 2.2 in [14]. Its Lipschitz constant is $\hat{L}(f)=1$. Using Proposition 8.1 one can derive a lower bound on the limiting learning of AdaGrad Norm.
@@
12 Numerical simulation details
@@
Concentration learning rate and risk for AdaGrad-Norm on a least squares problem with label noise $\omega=1$ (left) and on a logistic regression problem with no label noise (right). For logistic, see Section 11. 30 runs of AdaGrad-Norm with parameters $b=1$ and $\eta=1$ for each $d$; $X^{\star}\sim\mathcal{N}(0,I_{d}/d)$, $X_{0}=0$, and $K=I_{d}$. The shaded region represents a 90% confidence interval for the SGD runs. As the dimension increases, the risk and stepsize both concentrate around a deterministic limit (red). The deterministic limit is described by an ODE in Theorem 2.1. The initial loss increase in the least squares problem suggesting that the learning rate was initially too high, but AdaGrad-Norm naturally adapts and still the loss converges. Our ODEs predict this behavior.
@@
Quantities effecting AdaGrad-Norm learning rate. (left): The effect of adding noise to the targets ($\omega=1.0$) to the risk (left axis) and learning rate (right axis). Ran AdaGrad-Norm($b=1.0,\eta=2.5$) on least squares problem with $d=500$. $X_{0}$, $X^{\star}\sim\mathcal{N}(0,I_{d}/d)$. A single run of the SGD (solid line purple) matches exactly the prediction (ODE, teal). The shaded region represents $10$ runs of SGD with 90% confidence interval. The learning rate decays at the exact predicted rate of $\tfrac{\eta}{\sqrt{b^{2}+\frac{\text{Tr}{K}\omega^{2}}{d}t}}$. Depicted is learning $\tfrac{\text{learning rate}}{\text{asymptotic}}$ so it approaches $1$. (center, right): Noiseless least squares setting $(\omega=0)$. (center): Prop. 4.2 predicts the avg. eig of $K$ ($\text{Tr}(K)/d$) as compared with $\lambda_{\max}$ effects the $\lim_{k\to\infty}\mathfrak{g}_{k}$. Indeed, this is true. We varied the $\kappa=\lambda_{\max}/\lambda_{\min}$ while keeping the $\text{Tr}(K)/d$ and all other parameters fixed. All the learning rates behave identically verifying our theory about the effect of $\text{Tr}(K)/d$ on learning rates. (right): Varying the learning rate of AdaGrad norm by $\|X_{0}-X^{\star}\|^{2}$; our predictions (dashed) match and we see the inverse relationship predicted by Prop. 4.2. See Appendix 8 for details. Additionally, we did the following.
@@
Convergence in Exact Line Search on a noiseless least squares problem. The plot on the left illustrates the convergence of the risk function, while the center and right plots depict the convergence of the quotient $\frac{\mathscr{D}_{\lambda_{2}}(t)}{\mathscr{D}_{\lambda_{1}}(t)}$ and the learning rate $\gamma_{t}$, respectively. Predictions from ODE theory are compared with results obtained from SGD, demonstrating close agreement between the two approaches. Initialization was performed randomly, with $X_{0}\sim\mathcal{N}(0,I_{d}/\sqrt{d})$ and $X^{\star}\sim\tfrac{1}{\sqrt{d}}\mathbf{1}$, where $d=400$. The covariance matrix $K$ has two distinct eigenvalues $\lambda_{1}=1>\lambda_{2}>0$, and was constructed by specifying the spectrum, with $\lambda_{i}$ sampled from a discrete uniform distribution $\mathcal{U}\{1,\lambda_{2}\}$ for $i=1,\ldots,d=400$, and setting $K=\text{diag}(\lambda_{i}:i=1,\ldots,400)$. Further details and formulas for the limiting behavior can be found in the Appendix 10.2.
Finally, we include an example that uses real-world data, that is, the CIFAR 5m dataset [37]. Our theory does not explicitly deal with non-Gaussian distributions, but we find that the theoretical risk curves generalize cleanly to that case (the lighter coloured, stochastic paths represent the risk, whereas the darker, superimposed paths are the theoretical predictions).
Figure 6: Predict the training dynamics on a real dataset, CIFAR-5m [37], using multi-pass AdaGrad-Norm. This suggests the theory extends beyond Gaussian data and one-pass.
@@
Convergence on CIFAR 5m [37]. We train a classifier to distinguish between images of airplanes and cars. For a subset of the dataset with $n$ greyscale $32\times 32$ images, that is, $A\in\mathbb{R}^{n\times 1024}$, we randomly choose a matrix $W\in\mathbb{R}^{1024\times 2000}$ with i.i.d. Gaussian entries to generate the features $F=\text{relu}(AW)$. We then use multipass SGD with AdaGrad-Norm to find the least-squares mapping between $F$ and the labels $\ell$. All quantities are normalized to be constant order in $n$ and $d$. Finally, we compare with the ODE, provisionally modelling all the data as Gaussian. To run the ODE, we need two pieces of information: the eigenvalues of the covariance matrix, $F^{\top}F$, and the “noise." We define the noise as the least squares error: that is, the “variance" of the “noise" is
$\operatorname{\mathbb{E}}[\epsilon^{2}]=\min_{X\in\mathbb{R}^{d}}(FX-\ell)^{2}.$ (112)
We then run AdaGrad-Norm with initialization $X_{0}\sim d^{-1/2}\mathcal{N}(0,I_{d})$, $b=0.1$ and $\eta=10$, randomly sampling a datapoint from $F$ at every step. Note that the curves look significantly different for different $n$; smaller values of $n$ lead to an overparametrized problem, allowing least squares to memorize datapoints, whereas for larger $n$, least squares must learn a general function mapping images of cars and airplanes to their respective labels.
Acknowledgements
C. Paquette is a Canadian Institute for Advanced Research (CIFAR) AI chair, Quebec AI Institute (MILA) and C. Paquette was supported by a Sloan Fellowship, Google Grant (MILA), Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada, NSERC CREATE grant Interdisciplinary Math and Artificial Intelligence Program (INTER-MATH-AI), and FRQNT New University Researcher’s Start-Up Program; Research by E. Paquette was supported by a Google Grant (MILA) and Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada.

### v2

@@ D.2 Least squares setting
In the following we consider three cases, a strongly convex risk in which the spectrum of the eigenvalues is bounded from below (section D.2.1). A case in which the spectrum is not bounded from below as $d\to\infty$, but the number of eigenvalues below some fixed threshold is $o(d)$ (section D.2.2). Finally, power law spectrum supported on $[0,1]$ with $d\to\infty$ (section D.2.3).
D.2.1 Proofs for case of fixed $d$
@@
On the hand following Lemma D.3, $\frac{1}{4}\mathscr{D}^{2}(0)(1+\frac{\gamma_{0}}{2d}\operatorname{Tr}(K))\leq\|r\|_{1}$. Therefore, $\|r\|_{1}\asymp\frac{1}{4}\mathscr{D}^{2}(0)$.
Next, rewriting the $\gamma_{t}$ equation in Table LABEL:table:learning_rate_expressions (AdaGrad-Norm for least squares) in terms of $g(u)$ (Lemma D.1), we obtain
@@
Remark D.1.
We note that, on the Least square problem $\hat{L}(f)=\hat{\mu}(f)=1$, therefore, the bound in Proposition D.1 yields $\frac{\eta^{2}}{\frac{b}{\eta}+\frac{1}{2(1-\zeta)}\frac{1}{d}\operatorname{Tr}(K)\mathscr{D}^{2}(0)}$.
@@
D.2.2 Vanishingly few eigenvalues near 0 as $d\to\infty$
@@
with $\mathscr{R}_{2}(s)={\frac{1}{2d}}\sum_{i:\lambda_{i}\leq\alpha_{s}}\lambda_{i}\mathscr{D}_{i}^{2}(s)$. Next, choosing $\alpha_{s}=\gamma_{s}\frac{1}{d}\operatorname{Tr}(K^{2})$, we show that the last term is of order $o_{d}(1)$. By Lemma D.2 $\forall i$, $\mathscr{D}^{2}_{i}(t)\leq\max\left({\gamma_{t_{1}}\mathscr{R}(t_{1})},\mathscr{D}_{i}^{2}(0)\right)={c_{0}}$ where the bound ${c_{0}}$ comes from the assumption $\langle{X^{\star},\omega_{i}}\rangle=O(d^{-1/2})$ and the initialization $X_{0}=0$. Therefore,
@@
D.2.3 Power law behavior at $d\to\infty$
@@
In this section, we use the convolutional Volterra structure of the risk (Eq. (69)) to derive non-asymptotic bounds on the risk, which will be useful in Section D.2.3 to derive the asymptotic behavior of the risk and the learning rate under power law assumption on the spectrum of the covariance matrix and the discrepancy from the target at initialization.
Lemma D.3.
@@
where the third transition is by Lemma D.4, with $C_{1}=\frac{\mathcal{K}(0)}{\mathcal{K}(T)(2\epsilon+1)}+1,$ which then completes the proof. ∎
Lemma D.4 (Lemma IV.4.7 in [3]).
@@
Corollary D.1 apply Lemma D.3 in the setting for which $F$, and $\mathcal{K}$ has a power law behavior asymptotically. It shows that the risk will then be dominated by $F$ only. Corollary D.2 shows the behavior of the learning rate in this setting. Finally, Lemma D.5 shows that Assumption 8 is a consequence of a power law spectrum near zero on the eigenvalues of the covariance matrix and a power law assumption on the projected discrepancy at initialization.
Corollary D.1.
@@
Define $g(u)=\gamma_{\Gamma^{-1}(u)}$ and $r(u)=\mathscr{R}(\Gamma^{-1}(u))$ and observe that $g(u)$ is a decreasing function. Then, from the upper bound in Lemma D.3, we have
@@
Combining this upper bound with the lower bound from Lemma D.3 and that $\kappa_{2}>1$, we conclude that $r(u)\asymp F(u)$ and $\mathscr{R}(t)\asymp F(\Gamma(t)).$
@@
Using the notations $g(u)$ and $r(u)$ defined above along with the change of variable $u=\Gamma(t)$, we get $\int_{0}^{t}\mathscr{R}(s)\mathop{}\!\mathrm{d}s=\int_{0}^{u}\frac{r(v)}{g(v)}\mathop{}\!\mathrm{d}v$. Combining this with Corollary D.1 and the formula for $\gamma_{t}$ we get
@@
In the case of $\kappa_{1}<1$, we have $\sqrt{I(u)}-\sqrt{I(0)}\asymp(1+v)^{1-\kappa_{1}}$ so, for $u$ sufficiently large, $g(u)\asymp(1+u)^{\kappa_{1}-1}$. To recover the asymptotic for $\gamma_{t}$, we observe that $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}u}\Gamma^{-1}(u)=\frac{1}{g(u)}\asymp(1+u)^{1-\kappa_{1}}$. Integrating both sides and changing back to $t$ variables, we get $t\asymp(1+\Gamma(t))^{2-\kappa_{1}}$ (or equivalently $1+\Gamma(t)\asymp t^{1/(2-\kappa_{1})}$). Finally, plugging this into the formula for $\gamma_{t}$ and applying Corollary D.1, we get
@@
In the case of $\kappa_{1}=1$, we follow a similar procedure as for $\kappa_{1}<1$ to show that $t\asymp\Gamma(t)\log(\Gamma(t))$ for sufficiently large $t$. This implies $\Gamma(t)\asymp t/\log(t)$ which gives the desired result after integration. The decay rate of the risk is then immediate using Corollary D.1. ∎
Lemma D.5.
@@
The proof is an immediate application of Corollary D.2 with, $\kappa_{1}=2-\beta-\delta$ as implied by Lemma D.5. ∎
Remark D.2.
@@
Appendix E Polyak Stepsize
The distance to optimality of SGD is measured say by $D^{2}(X)=\|X-X^{\star}\|^{2}$. Let us consider the deterministic equivalent for the distance to optimality $\mathscr{D}^{2}(t)$ in (11). Fixing $T>0$ and any $\varepsilon\in(0,1/2)$, we have by Theorem 2.1 (see also corollary B.1 which show concentration for large class of statistics) that $\sup_{0\leq t\leq T}|\|X_{\lfloor td\rfloor}-X^{\star}\|^{2}-\mathscr{D}^{2}(t)|\leq d^{-\varepsilon}$, w.o.p. In this way, if we want to guarantee that the distance to optimality of SGD decreases, we need $\mathop{}\!\mathrm{d}\mathscr{D}^{2}(t)<0$ with the maximum decrease being $\min_{\gamma_{t}}\mathop{}\!\mathrm{d}\mathscr{D}^{2}(t)$.
@@ Appendix E Polyak Stepsize
Figure 5: Convergence in Exact Line Search on a noiseless least squares problem. The plot on the left illustrates the convergence of the risk function, while the center and right plots depict the convergence of the quotient $\frac{\mathscr{D}_{\lambda_{2}}(t)}{\mathscr{D}_{\lambda_{1}}(t)}$ and the learning rate $\gamma_{t}$, respectively. Further details and formulas for the limiting behavior can be found in the Appendix F.2. See Appendix H for simulation details.
Appendix F Line Search
F.1 General Line Search
@@
F.2 Line Search on least squares
@@ Appendix G Examples
G.1 Binary logistic regression
@@
We note that, the logistic regression is $(\mu,\theta)$–RSI with $\mu=\tfrac{1}{\ell e^{\sqrt{4\theta}}}$ see section 2.2 in [15]. Its Lipschitz constant is $\hat{L}(f)=1$. Using Proposition D.1 one can derive a lower bound on the limiting learning of AdaGrad Norm.
@@
G.2 CIFAR 5m
Finally, we include an example that uses real-world data, that is, the CIFAR 5m dataset [38]. Our theory does not explicitly deal with non-Gaussian distributions, but we find that the theoretical risk curves generalize cleanly to that case.
As we are now working with discrete data points rather than a distribution, the learning setup, while closely analogous to what was presented earlier, has some slight differences.
We start with a subset of the data consisting of $n$ grayscale images, each of which is $32\times 32$ pixels, that is, $A\in\mathbb{R}^{n\times 1024}.$ We fill a vector $b\in\mathbb{R}^{n}$ with the corresponding labels (00 for an image of a plane, $1$ for an image of a car.) We then randomly choose a matrix $W\in\mathbb{R}^{1024\times d}$ with i.i.d. Gaussian entries to generate the features $F=\text{relu}(AW)$. We want to use least squares to predict the label from the features, i.e., find
$\text{arg\,min}_{X\in\mathbb{R}^{d}}\left\{\mathcal{R}(X):=\frac{1}{2n}\left\|FX-b\right\|^{2}=\frac{1}{2n}\sum_{i=1}^{n}\left(f_{i}\cdot X-b_{i}\right)^{2}\right\},$ (111)
where $f_{i}$ is the $i$th row of $F$. The SGD we now consider is
$X_{k+1}=X_{k}-\gamma_{k}\left(f_{i_{k+1}}\cdot X-b_{i_{k+1}}\right)f_{i_{k+1}},\quad\left\{i_{k}\right\}\text{ iid }\operatorname{Unif}(\{1,2,\cdots,n\}),$ (112)
where $\gamma_{k}$ is the usual AdaGrad-Norm stepsize, as in (15). Our empirical covariance matrix $K$ (remembering that $f_{i}$ is a row vector) is then
$K={\mathbb{E}}\,_{i\in[n],j\in[n]}\left[f_{i}^{\top}f_{j}\right]=\frac{1}{n}F^{\top}F.$ (113)
[…]
Figure 6: Predicting the training dynamics on a real dataset, CIFAR-5m [38], using multi-pass AdaGrad-Norm. This suggests the theory extends beyond Gaussian data and one-pass. Note that the curves look significantly different for different $n$; smaller values of $n$ lead to an overparametrized problem, allowing least squares to memorize datapoints, whereas for larger $n$, least squares must learn a general function mapping images of cars and airplanes to their respective labels.
Appendix H Numerical simulation details
@@
Concentration learning rate and risk for AdaGrad-Norm on a least squares problem with label noise $\omega=1$ (left) and on a logistic regression problem with no label noise (right). For logistic, see Section G. 30 runs of AdaGrad-Norm with parameters $b=1$ and $\eta=1$ for each $d$; $X^{\star}\sim\mathcal{N}(0,I_{d}/d)$, $X_{0}=0$, and $K=I_{d}$. The shaded region represents a 90% confidence interval for the SGD runs. As the dimension increases, the risk and stepsize both concentrate around a deterministic limit (red). The deterministic limit is described by an ODE in Theorem 2.1. The initial loss increase in the least squares problem suggesting that the learning rate was initially too high, but AdaGrad-Norm naturally adapts and still the loss converges. Our ODEs predict this behavior.
@@
Quantities effecting AdaGrad-Norm learning rate. (left): The effect of adding noise to the targets ($\omega=1.0$) to the risk (left axis) and learning rate (right axis). Ran AdaGrad-Norm($b=1.0,\eta=2.5$) on least squares problem with $d=500$. $X_{0}$, $X^{\star}\sim\mathcal{N}(0,I_{d}/d)$. A single run of the SGD (solid line purple) matches exactly the prediction (ODE, teal). The shaded region represents $10$ runs of SGD with 90% confidence interval. The learning rate decays at the exact predicted rate of $\tfrac{\eta}{\sqrt{b^{2}+\frac{\text{Tr}{K}\omega^{2}}{d}t}}$. Depicted is learning $\tfrac{\text{learning rate}}{\text{asymptotic}}$ so it approaches $1$. (center, right): Noiseless least squares setting $(\omega=0)$. (center): Prop. 4.2 predicts the avg. eig of $K$ ($\text{Tr}(K)/d$) as compared with $\lambda_{\max}$ effects the $\lim_{k\to\infty}\mathfrak{g}_{k}$. Indeed, this is true. We varied the $\kappa=\lambda_{\max}/\lambda_{\min}$ while keeping the $\text{Tr}(K)/d$ and all other parameters fixed. All the learning rates behave identically verifying our theory about the effect of $\text{Tr}(K)/d$ on learning rates. (right): Varying the learning rate of AdaGrad norm by $\|X_{0}-X^{\star}\|^{2}$; our predictions (dashed) match and we see the inverse relationship predicted by Prop. 4.2. See Appendix D for details. Additionally, we did the following.
@@
Convergence in Exact Line Search on a noiseless least squares problem. The plot on the left illustrates the convergence of the risk function, while the center and right plots depict the convergence of the quotient $\frac{\mathscr{D}_{\lambda_{2}}(t)}{\mathscr{D}_{\lambda_{1}}(t)}$ and the learning rate $\gamma_{t}$, respectively. Predictions from ODE theory are compared with results obtained from SGD, demonstrating close agreement between the two approaches. Initialization was performed randomly, with $X_{0}\sim\mathcal{N}(0,I_{d}/d)$ and $X^{\star}\sim\tfrac{1}{\sqrt{d}}\mathbf{1}$, where $d=400$. The covariance matrix $K$ has two distinct eigenvalues $\lambda_{1}=1>\lambda_{2}>0$, and was constructed by specifying the spectrum, with $\lambda_{i}$ sampled from a discrete uniform distribution $\mathcal{U}\{1,\lambda_{2}\}$ for $i=1,\ldots,d=400$, and setting $K=\text{diag}(\lambda_{i}:i=1,\ldots,400)$. Further details and formulas for the limiting behavior can be found in the Appendix F.2.
@@
Convergence on CIFAR 5m [38]. We train a classifier to distinguish between images of airplanes and cars. Fix $d=2000$. Then for multiple values of $n$, we run AdaGrad-Norm with initialization $X_{0}=0$, $b=0.1$ and $\eta=5$, randomly sampling a datapoint from $F$ at every step. Details of the setup can be found in Appendix G.2.